# Towards the Automated Extraction and Refactoring of NoSQL Schemas from Application Code

### v2

@@ 5 Obtaining the Database Schema
Table 1 shows the DOS-to-U-Schema mappings applied to obtain logical schemas. The transformation begins by creating a USchemaModel as the root element. Then, an EntityType is created for each Container element. Next, each DataStructure is mapped to a StructuralVariation of the U-Schema metamodel, which includes a set of features. An entity type may have multiple variations if the code analysis is performed on different versions of the same script or application. Field elements are mapped to U-Schema features as follows: (i) An Attribute feature is created for each Attribute field, preserving its name and primitive type. (ii) A Reference feature is created for each Reference field, which is linked to the EntityType corresponding to the Container specified by the targetContainer property. (iii) Each Composition field is mapped to an embedded (non-root) entity type and an associated Aggregate feature. This process also creates a StructuralVariation for the referenced DataStructure, and is recursively applied to all fields contained within it. (iv) Each Collection field is mapped to an Attribute feature. The collection type is obtained from the collectionType property of the Collection meta-class. Collections may contain either primitive values or embedded objects.
Figure 13 shows the U-Schema model derived from the DOS model presented in Figure 11. The schema is visualized using SkiQL, a notation specifically designed for NoSQL schemas [10]. The schema includes the User and Movie entity types (yellow boxes), which correspond to the user and movie Containers. Each entity type aggregates (black dashed arrows) a single StructuralVariation. These variations are derived from the DataStructure associated with each container in the DOS model. A third entity type, WatchedMovie (grey box), represents the embedded data structure within the one associated with the user container. This entity type also includes a single structural variation, which defines a reference to Movie (blue solid arrow). An aggregate relationship (red solid arrow with a diamond) connects User to WatchedMovie. Attribute features are displayed inside the boxes representing the structural variations.
@@ 11       $node\leftarrow findFollowingNode(cfModel,join)$
12       while $\exists$ $node$ do
13             if $node.variables\in joinSList$land $node.variables\in prevSList$ then
@@ 29            $node\leftarrow getFollowingNode(node)$
Code is updated through a two-step process. First, a model-to-model transformation is applied to update the Code model. This transformation takes both the DOS and Code models as input and produces a modified Code model as output. The DOS model provides the necessary information: the data to be duplicated for each join query and the Read operations that can be removed. Accordingly, the identified join queries are eliminated, and the corresponding code expressions where the result variable was used are replaced. Specifically, the CodeBlock from the second query is moved into the first one. Then, all occurrences of the original result variable are replaced by the result variable of the preceding query, followed by access to the newly duplicated field. If the duplicated field belongs to an aggregate, the replacement must include access through the corresponding Aggregate field, as illustrated in the running example.
It is convenient to note that both deletion and replacement operations can be performed directly on the Code model, since it is cross-referenced by the database operations defined in the DOS model. In the second step, the updated Code model is traversed to regenerate the source code. This traversal had already been implemented as part of the testing process for Code model injection, as described in Section 3.
@@
Testing
To test the correctness of the code rewriting process, we compiled the transformed code using a standard JavaScript compiler and executed it against the modified database produced by the Orion engine. This ensured that the generated code was both syntactically valid and executable, and that the resulting queries produced the expected outputs when run on the refactored database.
@@ 7 Validation
First, we describe the experimental design and the methodology followed for the evaluation. Then, we present the results, and finally, we discuss the limitations of our validation.
7.1 Experiment: Description and Methodology
The validation was carried out through a round-trip experiment. Since publicly accessible schema and data are rarely available for non-trivial open-source MongoDB applications, we had to generate our own test environment. To address this, we generated all components of the test environment. The database schema (a U-Schema model) was created using the Athena language [16]; the dataset was generated with the data generator Deimos [15]. Both Athena and Deimos are tools specifically designed for working with U-Schema models. The application code was produced using an LLM, which was provided with the schema definition in Athena notation. The generated code simulated the backend logic of a small music streaming service, and is available on the git repository alongside the complete implementation of the approach.
Having the schema defined in advance facilitated the verification of the code analysis results. We were able to compare the U-Schema model obtained through our approach with both the original schema defined in Athena and the one inferred from the data using the approach presented in [11].
@@ 7.1 Experiment: Description and Methodology
Based on the schema shown in Figure 14, we generated a Node.js-based backend with the assistance of ChatGPT-4o, an LLM developed by OpenAI. The backend includes a REST API built with Express.js, with data stored in MongoDB and accessed via MongoDB native API [^10] MongoDB API: https://docs.mongodb.com/drivers/node. Each endpoint of the REST API performs at least one database operation. The domain model defines a class for each entity type in the schema and incorporates various structural relationships, including one-to-many and many-to-many references, as well as embedded documents. The codebase features realistic query patterns involving nested dereferencing and data aggregation, which were used to validate schema inference and the identification and application of refactorings such as join removal. In particular, several methods simulate join-like behavior by executing multiple dependent queries to resolve references across collections.
To populate the database, we created the collections artist, album, track, and genre to store the corresponding documents. Using our synthetic dataset generator, Deimos, we generated a representative and sufficiently diverse set of instances per entity type: 10 artists, 20 albums, 60 tracks, 8 genres, and 60 ratings. This configuration ensures a representative level of variability across relationships, such as tracks with multiple genres, albums and tracks with or without ratings, and nested access patterns through artist and album references. Although the dataset is limited in size, this does not affect the validity or generalizability of our results. Our approach is based on static code analysis, and therefore operates independently of the volume of data stored in the database. What matters for schema extraction and refactoring detection is the structure and usage of the application code—not the number of records. Once the relevant code access patterns are present, the size of the dataset has no influence on the models generated. This makes our validation both efficient and reproducible, without compromising its effectiveness in demonstrating the correctness of the analysis pipeline.
Next, we ran our code analysis solution using the generated application code as input. We first compared the inferred U-Schema model with the schema originally defined, and then with the schema inferred from the dataset using our data-based approach presented in [11]. These comparisons allowed us to assess the accuracy of our solution and evaluate how it compares with schema inference techniques based solely on stored data.
@@ 7.2 Results
# Query Description Source Entity Target Entity Field to Duplicate Duplication Location Join Type
@@ 7.3 Limitations of the validation
Finally, since the codebase used for validation was automatically generated, it may not fully reflect the diversity of patterns and idioms present in real-world applications. Future work should explore the application of this approach to existing systems with richer development histories and evolving schemas.
@@ 8 Related Work
Schema inference from graph stores is addressed in [8], where a model-driven reverse engineering approach is applied to analyze CREATE statements written in Neo4j’s Cypher query language. This analysis enables the extraction of node types, relationships, and their properties, which are then represented in a graph-based metamodel. The resulting model is subsequently transformed into an Extended Entity-Relationship (EER) schema for conceptual modeling purposes.
@@
U-Schema was proposed by [11] as a unified metamodel capable of representing logical schemas for both relational and NoSQL databases, including document, key–value, columnar, and graph models. The authors defined canonical forward and reverse mappings between U-Schema and each supported data model. A common strategy was established for implementing and validating the schema extraction process across all five types of databases. For validation, synthetic data were generated to populate the databases, and a four-step round-trip experiment was conducted. The main advantages of the U-Schema-based approach include its independence from any specific data model and its ability to extract schemas that capture both reference and aggregation relationships, as well as structural variations. Moreover, the resulting schema is a model that conforms to an Ecore/EMF metamodel [25], which enables the use of EMF tooling to build database utilities and facilitates interoperability with other tools, as discussed in [22].
As indicated in Section 1, our code analysis approach is complementary to the data-driven strategy proposed in [11]. While analyzing application code allows reference relationships to be directly identified—without relying on heuristics, which may not always be reliable—structural variability is more effectively discovered through data analysis. In contrast, detecting schema variations through code requires access to multiple versions of the application over time. The other advantages discussed above —such as being a generic approach, extracting schemas with references and aggregation, and producing models conforming to Ecore— also apply to our code analysis solution.
@@
Several metamodels have been proposed to represent the structure and behavior of source code in a language-independent manner. These metamodels are widely used in static analysis, reverse engineering, and software modernization tasks, as they provide an abstract representation of code elements such as classes, methods, variables, and control flow. In the following, we briefly describe two of the most relevant metamodels —KDM and MoDisco— which have influenced the design of our Code metamodel and are widely recognized in the context of model-driven reverse engineering.
KDM is a specification defined by the Object Management Group [21]. It is a comprehensive metamodel composed of multiple packages that enable the representation of various aspects of software systems, ranging from source code to physical deployment. KDM is designed to support a wide variety of programming languages, and it provides support to be extended in order to capture language-specific constructs. While our Code metamodel was inspired by KDM, it is intentionally kept simpler and focused on representing core object-oriented constructs that are sufficient for analyzing applications with intensive database interaction.
@@ 9 Conclusions and Future Work
Finally, the code metamodel could be extended or evolved into a family of metamodels to support different programming paradigms. An alternative and potentially more scalable strategy would be to define a modular set of metamodels, each tailored to a specific paradigm. A significant challenge in this direction is the representation of multi-paradigm languages—such as Scala or Rust—which blend object-oriented, functional, and imperative features within a single codebase.
@@ Acknowledgements
The contributions of the first and third authors were supported by project PID2020-117391GB-I00, funded by MICIU/AEI /10.13039/501100011033 (Spain), and co-funded by ERDF/EU; Anthony Cleve was supported by the Fonds de la Recherche Scientifique (F.R.S.-FNRS) under the PDR project INSTINCT (35270712).

### v3

@@ 5 Obtaining the Database Schema
The transformation begins by creating a USchema-Model as the root element. Then, an EntityType is created for each Container element. Next, each DataStructure is mapped to a StructuralVariation of the U-Schema metamodel, which includes a set of features. An entity type may have multiple variations if the code analysis is performed on different versions of the same script or application. Field elements are mapped to U-Schema features as follows: (i) An Attribute feature is created for each Attribute field, preserving its name and primitive type. (ii) A Reference feature is created for each Reference field, which is linked to the EntityType corresponding to the Container specified by the targetContainer property. (iii) Each Composition field is mapped to an embedded (non-root) entity type and an associated Aggregate feature. This process also creates a StructuralVariation for the referenced DataStructure, and is recursively applied to all fields contained within it. (iv) Each Collection field is mapped to an Attribute feature. The collection type is obtained from the collectionType property of the Collection meta-class. Collections may contain either primitive values or embedded objects.
Table 2 shows the DOS to U-Schema mappings applied to obtain logical schemas. Figure 13 shows the U-Schema model derived from the DOS model presented in Figure 11. The schema is visualized using SkiQL, a notation specifically designed for NoSQL schemas [11].
@@
Testing
As explained in the previous sections, each transformation step was validated independently. Manual techniques were applied for the Code and DOS models, using incremental testing with code snippets of increasing complexity, and for the Control Flow models, through visual inspection of the generated graphs. In contrast, the final transformation (from DOS to U-Schema) was automatically validated using Athena [17], a generic schema definition language built on top of U-Schema. Athena enabled automated testing by generating and comparing reference scripts, providing an objective mechanism to verify the correctness of the transformation and, consequently, of the entire workflow.
@@ 11  $node\leftarrow findFollowingNode(cfModel,join)$
12  while $\exists$ nodenode do
13     if $node.variables\in joinSList$ $\land$ $node.variables\in prevSList$ then
@@ 29    $node\leftarrow getFollowingNode(node)$
Code is updated through a two-step process. First, a model-to-model transformation is applied to update the Code model. This automatic transformation takes both the DOS and Code models as input and produces a modified Code model as output. The DOS model provides the necessary information: the data to be duplicated for each join query and the Read operations that can be removed. Accordingly, the identified join queries are eliminated, and the corresponding code expressions where the result variable was used are replaced. Specifically, the CodeBlock from the second query is moved into the first one. Then, all occurrences of the original result variable are replaced by the result variable of the preceding query, followed by access to the newly duplicated field. If the duplicated field belongs to an aggregate, the replacement must include access through the corresponding Aggregate field, as illustrated in the running example.
Finally, in the second step, the updated Code model is traversed to regenerate the source code. Note that the regenerated code is automatically modified and may contain minor issues that require manual review.
@@ 7 Validation
First, we describe the experimental design and the me-thodology followed for the evaluation. Then, we present the results, and finally, we discuss the limitations of our validation.
7.1 Experimental setup
Because no publicly available datasets provide both the source code and its corresponding database schema as a reliable ground truth, we generated our own validation environment. In practice, the intended logical schema of real applications is seldom documented or formally specified, which prevents an objective assessment of the accuracy of the inferred models.
To archieve this, we generated all components of the test environment. The database schema (U-Schema model) was created using the Athena language [17]; the dataset was generated with the data generator Deimos [16]. Both Athena and Deimos are tools specifically designed for working with U-Schema models. The application code was automatically generated using a large language model (LLM), which was provided with the schema definition. The resulting code simulated the backend logic of a small music streaming service, following common JavaScript development statements. This strategy allowed us to create diverse and reproducible examples where the intended schema was known in advance, thus enabling a systematic validation of our approach despite the scarcity of real-world datasets. The generated project and the complete implementation of the proposed approach are publicly available in the GitHub repository referenced in Section 1.
@@ 7.1 Experimental setup
Based on the schema shown in Figure 14, we generated a Node.js-based backend with the assistance of ChatGPT-4o, an LLM developed by OpenAI. The model was provided with the schema as input (in the form of Figure 14) and was instructed to generate a fully functional application using the MongoDB native API driver in JavaScript, while choosing the implementation details it deemed most appropriate. Before code generation, we asked the model clarifying questions to verify that it had correctly understood all entities, attributes, and references, including their names and relationships. The resulting backend includes a REST API built with Express.js, with data stored in MongoDB and accessed via the MongoDB native API [^8] MongoDB API: https://docs.mongodb.com/drivers/node. The generated code included CRUD operations for all entities defined in the schema and corresponding endpoint handlers. We manually verified that all schema elements were correctly represented and handled in the generated code.
Each endpoint of the REST API performs at least one database operation. The domain model defines a class for each entity type in the schema and incorporates various structural relationships, including one-to-many and many-to-many references, as well as embedded documents. The codebase features realistic query patterns involving nested dereferencing and data aggregation, which were used to validate schema inference and the identification and application of refactorings such as join removal. In particular, several methods simulate join-like behavior by executing multiple dependent queries to resolve references across collections.
To populate the database, we created the collections artist, album, track, and genre to store the corresponding documents. Using our synthetic dataset generator, Deimos, we generated a representative and sufficiently diverse set of instances per entity type: 10 artists, 20 albums, 60 tracks, 8 genres, and 60 ratings. This configuration ensures a representative level of variability across relationships, such as tracks with multiple genres, albums and tracks with or without ratings, and nested access patterns through artist and album references. Although the dataset is limited in size, this does not affect the validity or generalizability of our results. Our approach is based on static code analysis, and therefore operates independently of the volume of data stored in the database. What matters for schema extraction and refactoring detection is the structure and usage of the application code—not the number of records. Once the relevant code access patterns are present, the size of the dataset has no influence on the models generated.
7.2 Methodology: Round Trip experiment
The validation was carried out through a round-trip experiment. We used the application code generated from the previously defined database schema as input to our code analysis approach. The inferred U-Schema model resulting from our approach was then compared with the original schema and with the schema inferred from the dataset using our data-based approach presented in [12]. These comparisons allowed us to assess the accuracy of our solution and evaluate how it performs relative to our data-based approach.
@@ 7.3 Results
# Query Description Target Entity Source Entity Field to Duplicate Duplication Location Join Type
@@ 7.4 Limitations of the validation
It is important to note that the projects generated with the help of LLMs are more regular and simplified than typical JavaScript applications. Real-world projects often include heterogeneous coding practices, extensive use of asynchronous constructs, and callbacks where database objects are passed across different functions or modules. Such cases are challenging for static analysis, as the control flow may only be resolved at runtime, and are not fully represented in our current evaluation dataset. Nevertheless, the LLM-assisted experiment proved useful to identify the limits of our approach when contrasted with real-world development practices. It helped us better characterize the kinds of constructs that hinder static analysis—such as asynchronous callbacks, middleware chaining, dynamic typing, and indirect persistence layers that obscure database interactions. This limitation is explicitly acknowledged, and we outline in Section 8 potential directions to mitigate it, including hybrid static/dynamic analysis and the analysis of applications written in TypeScript.
@@ 8 Related Work
Schema inference from graph stores is addressed in [9], where a model-driven reverse engineering approach is applied to analyze CREATE statements written in Neo4j’s Cy-pher query language. This analysis enables the extraction of node types, relationships, and their properties, which are then represented in a graph-based metamodel. The resulting model is subsequently transformed into an Extended Entity-Relationship (EER) schema for conceptual modeling purposes.
@@
U-Schema was proposed by [12] as a unified metamodel capable of representing logical schemas for both relational and NoSQL databases, including document, key–value, columnar, and graph models. The authors defined canonical forward and reverse mappings between U-Schema and each supported data model. A common strategy was established for implementing and validating the schema extraction process across all five types of databases. For validation, synthetic data were generated to populate the databases, and a four-step round-trip experiment was conducted. This USchema-based approach offers signficant advantages over the approaches discussed above that include its independence from any specific data model and its ability to extract schemas that capture both reference and aggregation relationships, as well as structural variations. Moreover, the resulting schema is a model that conforms to an Ecore/EMF metamodel [26], which enables the use of EMF tooling to build database utilities and facilitates interoperability with other tools, as discussed in [23].
Building upon the same U-Schema foundation, our code analysis approach complements the data-driven strategy proposed in [12]. While analyzing application code allows reference relationships to be directly identified —without relying on heuristics, which may not always be reliable— structural variability is more effectively discovered through data analysis. In contrast, detecting schema variations through code requires access to multiple versions of the application over time. The other advantages discussed above —such as being a generic approach, extracting schemas with references and aggregation, and producing models conforming to Ecore— also apply to our code analysis solution.
@@
Two previous works from our group are also closely related to the present approach. Sánchez-Ramón et al. [22] proposed a reverse engineering strategy to recover the event-handling logic of Oracle Forms applications. Their work analyzed PL/SQL triggers and identified code fragments corresponding to interaction idioms, building a Control Flow model based on Ullmann’s algorithm [1]. The resulting model represented control dependencies between UI, control, and business-logic layers. Although the goal was not database schema extraction, this work shares with ours the use of static analysis and control-flow reconstruction to reveal implicit structures in legacy code.
Fernández-Candel et al. [10] presented a model-driven reengineering process for migrating PL/SQL triggers to Java organized into MVC layers. That work defined Code and Action models compliant with KDM and employed a testing strategy similar to the one used here to validate model transformations. The current proposal extends these ideas to a different domain—database access reconstruction—by defining a specialized Code metamodel and a chain of model transformations that culminate in the automated derivation of database schemas.
@@
Several metamodels have been proposed to represent the structure and behavior of source code in a language-independent manner. These metamodels are widely used in static analysis, reverse engineering, and software modernization tasks, as they provide an abstract representation of code elements such as classes, methods, variables, and control flow. In the following, we briefly describe two of the most relevant metamodels —KDM and MoDisco— which have influenced the design of our Code metamodel, as explained at the beginning of Section 3.1, and are widely recognized in the context of model-driven reverse engineering.
KDM is a specification defined by the Object Management Group [21]. It is a comprehensive metamodel composed of multiple packages that enable the representation of various aspects of software systems, ranging from source code to physical deployment. KDM is designed to support a wide variety of programming languages, and it provides support to be extended in order to capture language-specific constructs. While our Code metamodel was inspired by the language-independent design of KDM’s Code package, it was intentionally kept simpler and focused on representing the core object-oriented constructs required for analyzing database-intensive applications.
@@
Table 4 compares the approaches according to their analyzed artefacts (Input), produced artefacts (Output), analysis technique, intermediate representations, and their degree of language and database independence. The last column (Application) indicates how the inferred models are used or applied, when applicable. While data-driven approaches are inherently constrained by the coverage and representativeness of the available data—that is, they can only infer structures present in the analyzed instances—, code-based approaches face challenges with language dependence and dynamic constructs, and refactoring methods are often technology-specific. In contrast, our proposal uniquely combines static code analysis with dedicated metamodels and a language- and database-independent representation, enabling systematic schema inference and technology-agnostic refactoring.
Work Input Output Kind of analysis Intermediate representation Language/DB independence Application
Wang et al. (2015) [28] Database content JSON Schema Data-based Tree No/No (MongoDB) Schema inference
Klettke et al. (2015) [19] Database sample JSON Schema Data-based Graph No/No (MongoDB) Schema inference
Comyn-Wattiau et al. (2017) [9] CREATE statements EER schema Data-based Graph-based metamodel No/No (Neo4j) Schema inference
Frozza et al. (2021) [14] Database content JSON Schema Data-based No/No (HBase) Schema inference
Fernández-Candel et al. (2022) [12] Database content U-Schema model Data-based Metamodels for Relational and NoSQL paradigms Yes/Yes Schema inference
Cherry et al. (2022) [6] Source Code (JavaScript) Access map Static (code-driven) No/No (MongoDB) Query analysis
Bernard et al. (2021) [3] Source Code (Java) Code patterns Static (code-driven) No/No (MongoDB) Query analysis
Meurice et al. (2017) [20] Source Code (Java) Schema evolution model Static (code-driven) No/No (MongoDB) Database evolution
Hernández-Chillón et al. (2023) [8] Source Code (Java) Database Changes model Static (code-driven) Orion metamodel Yes/No (MongoDB) Schema–code co-evolution
Sánchez Ramón et al. (2011) [22] Event Handlers (PL/SQL) Java Code Static (code-driven) Behaviour and Event-Handler metamodels Yes/No (Oracle) Software Reverse Engineering
Fernández-Candel Carlos et al. [10] Source Code (PL/SQL) Java Code Static (code-driven) KDM (Code and Action) metamodel Yes/No (Oracle) Software Reverse Engineering
Our approach Source Code (Javascript) U-Schema model + Refactoring plan Static (code-driven) Code, Control Flow & DOS metamodels Yes/Yes Schema inference and refactoring
Table 4: Comparative analysis of existing approaches addressing database schema inference and source code analysis.
@@ 9 Conclusions and Future Work
One important limitation of our current static analysis is its handling of callbacks and asynchronous programming patterns, which are particularly frequent in JavaScript. Our algorithm can follow explicit calls in the same module, more complex situations—such as objects passed to external callbacks, asynchronous chains, or dynamically registered handlers—cannot be fully resolved statically, since the exact control flow may only be determined at runtime.
@@
Recent advances in large language models (LLMs) specialized in code understanding —such as OpenAI’s Codex, CodeLlama, or Anthropic’s Claude Sonnet 4.5— open pro-mising opportunities to complement static analysis. These models can reason about implicit semantics, incomplete code fragments, and dynamically generated constructs that are difficult to capture through purely syntactic or control-flow–based methods. However, their outputs are non-deter-ministic, context-dependent, and lack traceability, which complicates reproducibility and validation—key aspects in both research and industrial contexts. Moreover, LLMs require substantial computational resources and may introduce biases or inaccuracies when reasoning over complex or domain-specific codebases.
A promising research direction is to develop hybrid approaches that integrate LLM-based reasoning with static analysis. In such a combination, static analysis would provide systematic coverage and formal guarantees, while LLMs could contribute higher-level reasoning and pattern recognition capabilities, achieving a balance between reliability and flexibility. Hybrid static/dynamic solutions, such as workload monitoring or runtime tracing, could also improve the accuracy of schema inference in realistic settings where control flow depends on runtime conditions.
Our current validation is based on synthetic projects generated with the assistance of LLMs. This strategy allowed us to control the scenarios and obtain ground truth, but the resulting code was more regular than typical Java-Script applications. Real-world projects often exhibit heterogeneous coding practices, including extensive use of asynchronous patterns and callbacks where query results are passed across functions or modules. Such cases are only partially captured by our static analysis, since the exact control flow may only be determined at runtime. Finally, future work will extend the base metamodel by adding language-specific extensions. We plan to start with JavaScript and later experiment with Python and TypeScript, which will require implementing injectors from their ASTs to the Code model. Supporting multi-paradigm languages such as Scala would be considerably more complex, as it would involve modeling functional and hybrid constructs beyond the current object-oriented scope.
@@ Acknowledgements
The contributions of the first and third authors were supported by project PID2020-117391GB-I00, funded by MICIU/AEI /10.13039/501100011033 (Spain), and co-fun-ded by ERDF/EU; Anthony Cleve was supported by the Fonds de la Recherche Scientifique (F.R.S.-FNRS) under the PDR project INSTINCT (35270712).